# Confronting compositional confusion through the characterisation of the sub-Neptune orbiting HD 77946

### v1

@@ Abstract
We report on the detailed characterization of the HD 77946 planetary system. HD 77946 is an F5 ( = 1.17 M, = 1.31 R) star, which hosts a transiting planet recently discovered by NASA’s Transiting Exoplanet Survey Satellite (TESS), classified as TOI-1778 b. Using TESS photometry, high-resolution spectroscopic data from HARPS-N, and photometry from CHEOPS, we measure the radius and mass from the transit and RV observations, and find that the planet, HD 77946 b, orbits with period = d, has a mass of M, and a radius of R. From the combination of mass and radius measurements, and the stellar chemical composition, the planet properties suggest that HD 77946 b is a sub-Neptune with a 1% H/He atmosphere. However, a degeneracy still exists between water-world and silicate/iron-hydrogen models, and even though interior structure modelling of this planet favours a sub-Neptune with a H/He layer that makes up a significant fraction of its radius, a water-world composition cannot be ruled out, as with K, water may be in a supercritical state. The characterisation of HD 77946 b, adding to the small sample of well-characterised sub-Neptunes, is an important step forwards on our journey to understanding planetary formation and evolution pathways. Furthermore, HD 77946 b has one of the highest transmission spectroscopic metrics for small planets orbiting hot stars, thus transmission spectroscopy of this key planet could prove vital for constraining the compositional confusion that currently surrounds small exoplanets.
@@
The ‘radius valley’ that separates super-Earths and sub-Neptunes has been consistently observed at 1.5 - 2 R (Fulton et al., 2017; Van Eylen et al., 2018; Fulton & Petigura, 2018; Ho & Eylen, 2023), and is largely without planets. Debate currently surrounds the origin of this gap, with proposed scenarios including core-powered mass loss, photoevaporation, or that these planets are primordially rocky. Interpretations differ on the physical mechanism of atmospheric mass-loss, but the result is the same - primordially accreted atmospheres are removed in such a way that different planets are affected in different ways over different timescales, resulting in a valley which separates a population of stripped-core planets (super-Earths) from those which have retained their H/He envelopes (sub-Neptunes) (Owen & Wu, 2013; Ginzburg et al., 2018; Gupta & Schlichting, 2019). Whilst investigations have been launched into the existence of planets in the valley, it is still not immediately clear if there are any firm detections of planets in the gap. Larger population studies have indicated that at least a few planets fall within the radius valley, but conversely smaller and more precise planet samples have revealed a complete lack of planets across the gap (Ho & Eylen, 2023). It has been observed, however, that the valley varies with planetary orbital separation, system age, and mass and metallicity of the host stars, which is consistent with the predictions of atmospheric mass-loss models (David et al., 2021; Sandoval et al., 2021; Petigura et al., 2022).
The internal structure of sub-Neptunes is not just limited to that of a rocky core surrounded by a gaseous atmosphere, it has been theorised that these planets might hold significant fractions of ices or liquid water (Rogers & Seager, 2010; Dorn et al., 2017). Zeng et al. (2021) suggested that the radii of planets hotter than 900K and with masses below 20 M can be reproduced assuming ice-dominated compositions without significant gaseous envelopes. The entire mass-radius (M-R) distribution of planets around the radius valley could therefore be separated into three groups: purely rocky planets, gas-poor water worlds, and gas-rich water worlds, where the first two represent the two peaks in the radius distribution (Luque & Pallé, 2022). However, Rogers et al. (2023) argues that the existence of small planets with hydrogen atmospheres are consistent with the data, once thermal evolution and mass-loss are properly accounted for. This means that there is a strong degeneracy between water-world and silicate/iron-hydrogen models, and that the characterisation of larger sub-Neptunes that reside in this area of the M-R diagram can be used to determine planetary evolution and formation pathways (Kubyshkina & Fossati, 2022).
@@ 2.1.2 CHEOPS
The data were processed using the CHEOPS Data Reduction Pipeline (DRP v13; Hoyer et al. 2020) that conducts frame calibration, instrumental and environmental correction, and aperture photometry using pre-defined radii ( = 22.5″[RINF], 25.0″[DEFAULT], 30.0″[RSUP], and OPTIMAL radius) as well as a noise-optimised radius [ROPT]. The size of the OPTIMAL radius is determined on a per-visit basis to account for the differing brightnesses of different targets, and is determined by minimizing the signal-to-noise ratio (Hoyer et al., 2020). Both visits of HD 77946 have the same optimized aperture radius of 25 pixels. The DRP produced flux contamination (see Hoyer et al. 2020 and Wilson et al. 2022 for computation and usage) that was subtracted from the light curves. We retrieved the data and corresponding instrumental basis vectors, assessed the quality using the pycheops Python package (Maxted et al., 2022), and decorrelated with the parameters suggested by this package, which can be found in Table 1. Outliers were also trimmed from the light curves, with points that were 4 away from the median value removed. We concluded that using the OPTIMAL radius was the best option as it gave the highest signal-to-noise. Therefore, we used these detrended data for further analysis.
@@ 2.2.1 DRS
Following the confident detection of the transiting planet with the DRS RVs, the s-BART (Silva et al., 2022) template-matching pipeline was applied to the same data-set. This approach is built around the core assumption that the RV signal introduced by the presence of an orbiting planet should be achromatic, i.e. independent of the wavelength at which it is measured. Telluric features are accounted for through a binary mask that rejects any wavelength for which Earth’s transmittance drops under 1% of its continuum value, which is enlarged to consider the BERV motion. Furthermore, to ensure that each observation is informed by the same wavelength region, s-BART only uses the spectral orders that are common to all observations in the dataset. This model was implemented in a Bayesian framework, allowing for a straightforward and consistent characterization of the uncertainties associated with the RV extraction, and has been yielding good results for both ESPRESSO data (e.g. Faria et al. 2022) and HARPS-N (Wilson et al 2024, in review). This resulted in an average RV precision of 1.4 ms-1 and an RMS of 4.45 ms-1. The HARPS-N s-BART data are presented in Table 6.
@@ 3.1 Atmospheric parameters
We calculate the U, V, and W Galactic velocities using the method of Johnson & Soderblom (1987) and data from the most recent Gaia data release (for examples see Mortier et al. (2020); Wilson et al. (2022)), and compute the thin and thick disk, and halo membership probabilities. This was done using the framework outlined in Wilson et al. (2024) (in review), which combines new Gaia offset-corrected (Lindegren et al., 2021) data (Gaia Collaboration et al., 2022) with a known probability method (Reddy et al., 2006), then uses a novel Monte Carlo approach to compute a weighted average combination of probabilities using four sets of velocity standards (Bensby et al., 2003; Bensby et al., 2014; Reddy et al., 2006; Chen et al., 2021) corrected for the Local Standard of Rest (Koval et al., 2009). We find U kms-1, V kms-1, W kms-1, resulting in 98.5%, 1.5%, and 0% thin disk, thick disk, and halo probabilities, respectively, which is in line with the given metallicity for this type of star.
@@
| RA | 09:07:06.63 | Gaia EDR3 |
| Declination | +46:40:20.73 | Gaia EDR3 |
| [mas] | Gaia EDR3 | |
@@
| P [d] | This work | |
| RV [m s] | This work | |
| [km s] | This work | |
| [km s] | This work | |
| [km s] | This work | |
@@ 3.2 Stellar mass, radius, and age
Stellar parameters of HD 77946 were determined using isochrones and stellar evolution tracks. For a detailed discussion on this see Mortier et al. (2020). In brief, spectroscopically determined values for the effective temperature and metallicity from ARES+MOOG and SPC were used as inputs to the isochrones python package (Morton, 2015), together with the Gaia Early Data Release 3 (Brown et al., 2021) parallax and the apparent magnitude in 8 bands, listed in Table 2. Two sets of stellar evolution models were used, the Dartmouth Stellar Evolution Database (Dotter et al., 2008) and the MESA Isochrones and Stellar Tracks (MIST - Dotter, 2016). For each stellar evolution model (Dartmouth/MIST) the analysis was run twice; once using each of the spectroscopic inputs (ARES+MOOG/SPC), totalling to 4 individual runs. Final values for the stellar mass, radius, and age were obtained by combining the posteriors from each run, and taking the median and the 16th and 84th percentile. We derived values M, R, and Gyr. Combining the derived mass and radius, we can compute a new value for the surface gravity, which is consistent with, but more precise than the spectroscopic surface gravity.
@@ 3.3 Stellar activity and rotation
Signals intrinsic to stars themselves still pose significant problems, as they can induce RV variations that can drown out or even mimic planetary signals (Meunier, 2021). Modelling and thus mitigating these signals is notoriously difficult. However, a technique that has proved successful for disentangling stellar activity signals from planetary signals is to use a Gaussian Process (GP) framework (Haywood et al., 2014; Rajpaul et al., 2015; Stock et al., 2020). GPs have a number of features that make them very well suited to the joint modelling of activity processes and dynamical (e.g. planetary) signals, giving probabilistic estimates of the model parameters, which leads to more precise and accurate mass estimates.
@@
where and are two times of observation, and the other terms are the variability period of the star (), the variability amplitude (), the periodic characteristic length (associated with the number of spots/spot regions on the surface of the star, ), and the non-periodic characteristic length (associated with the spot decay timescale, ). The GP was run with log-uniform priors on the amplitude 0 ln() 2.7 and active-region evolution timescale 24 40 d. These boundaries were selected based on the stellar type of HD 77946, an F dwarf, so we assume its active regions evolve on approximately the same timescale as the Sun. The SPC gives kms, and based on a stellar radius of 1.31 R this limits the stellar rotation period to a value of below 22 days. Based on this estimate and a few initial runs with uninformative priors, the GP priors on the stellar rotation period were set to be uniform with 12 19 d. The log harmonic complexity was controlled with a uniform prior -1 ln() 0. In short, we used a model with up to five unknown Keplerian signals to maximise exploration, and stellar activity decorrelation performed against the scalpels basis vectors, incorporated with a quasi-periodic GP regression.
@@ 4.1 Joint Transit & RV Modelling
| [M] | Planet mass | |||
| [R] | Planet radius | |||
@@
Inferred stellar and planetary parameters from the combined fit are shown in Table 3, alongside the priors implemented. The prior for is explained in Section 3.3, whilst the prior on stellar density comes from stellar mass and radius estimates in Section 3.2. The priors placed on the limb-darkening coefficients were calculated using interpolation of Table 25 from Claret (2017) for the TESS light curves and Table 8 from Claret (2021) for the CHEOPS light curves. The fit converges to a moderately eccentric planetary orbit, with a posterior on corresponding to a detection. Using the estimates for the stellar mass and radius obtained in Section 3.2 and the HARPS-N DRS RVs, we derive that HD 77946 b has a mass M and a radius R.
We also performed the same combined GP, transit, and RV fit using the s-BART RVs, as mentioned in Section 2.2.2, with the resulting inferred stellar and planetary parameters shown in Table 3. In this case the fit converged to a slightly more eccentric planetary orbit than that of the fit with the DRS RVs, which led to a slightly higher posterior on corresponding to a detection. From this we also derived a slightly lower mass and radius for HD 77946 b of M and R. These results still fall within of those derived from the HARPS-N DRS RVs, but this does suggest that the choice of data reduction technique can influence the resulting planetary parameters. We favour these results as we find that the fit with the s-Bart RVs gives a stronger detection of the planet’s mass and required less jitter in the fit, although there is tension in the derived eccentricity when using either RV dataset.
@@
We ensured that the nested sampling was converging to consistent results by running the same model multiple times, and across different machines. Given the aforementioned problems in constraining the stellar rotation period, we also ran models with different priors on , despite this the planetary parameter results were still consistent to 1 on all runs. A corner plot with all the planetary parameters can be found in Appendix A, see Figure 11. In addition to this the RV semi-amplitude (2.69 ± 0.57 ms) obtained from the scalpels analysis in Section 3.3 was combined with the stellar mass and orbital inclination (1.17 M and 87.10 degrees respectively) which lead to a mass determination of 8.83 ± 1.86 M. This result is consistent with the results from the joint GP, photometric, and spectroscopic fit determined using PyORBIT. Fits were also performed with and without a GP, and comparing the Bayesian Information Criterion (BIC), the non-GP fit had a log-median BIC of -64134 whilst the fit with a GP had a log-median BIC of -56987. Trotta (2008) suggests under the assumption of a multivariate Gaussian posterior that a log-Bayes factor of should be considered "strong" evidence for a model. In this case the log-Bayes factor is 3573.5 (BIC), further justifying the use of a GP.
In order to further confirm this, a phase folded plot of the HARPS-N RVs colour coded to the time period they were taken was produced (see Appendix A Figure 10). From the grouping of RVs in this plot, we were able to conclude that the results were not being skewed by only sampling some parts of the orbital phase. An additional check was performed to ensure the validity of the eccentricity obtained, in which a 1-planet fit was attempted where the planets orbit was forced to be circular (with ). From this we calculate a slightly lower mass, M, corresponding to a similar confidence detection of , with R. Comparing the Bayesian Information Criterion (BIC), the forced circular orbit had a log-median BIC of -57009 whilst the eccentric orbit had a log-median BIC of -56987. Thus, whilst the confidence detections on are both around , the median BIC of the eccentric orbit is better. In addition to this, the log-Bayes factor is 11 (BIC), which supports choosing the more eccentric model. Orbits for single-planet systems are also not typically zero (Van Eylen et al., 2019) and there is no compelling argument to fix the eccentricity of this planet as circular, therefore we favour the model with an eccentric orbit.

### v2

@@ Abstract
We report on the detailed characterization of the HD 77946 planetary system. HD 77946 is an F5 ( = 1.17 M⊙, = 1.31 R⊙) star, which hosts a transiting planet recently discovered by NASA’s Transiting Exoplanet Survey Satellite (TESS), classified as TOI-1778 b. Using TESS photometry, high-resolution spectroscopic data from HARPS-N, and photometry from CHEOPS, we measure the radius and mass from the transit and RV observations, and find that the planet, HD 77946 b, orbits with period = d, has a mass of M⊕, and a radius of R⊕. From the combination of mass and radius measurements, and the stellar chemical composition, the planet properties suggest that HD 77946 b is a sub-Neptune with a 1% H/He atmosphere. However, a degeneracy still exists between water-world and silicate/iron-hydrogen models, and even though interior structure modelling of this planet favours a sub-Neptune with a H/He layer that makes up a significant fraction of its radius, a water-world composition cannot be ruled out, as with K, water may be in a supercritical state. The characterisation of HD 77946 b, adding to the small sample of well-characterised sub-Neptunes, is an important step forwards on our journey to understanding planetary formation and evolution pathways. Furthermore, HD 77946 b has one of the highest transmission spectroscopic metrics for small planets orbiting hot stars, thus transmission spectroscopy of this key planet could prove vital for constraining the compositional confusion that currently surrounds small exoplanets.
@@
The ‘radius valley’ that separates super-Earths and sub-Neptunes has been consistently observed at 1.5 - 2 R⊕ (Fulton et al., 2017; Van Eylen et al., 2018; Fulton & Petigura, 2018; Ho & Eylen, 2023), and is largely without planets. Debate currently surrounds the origin of this gap, with proposed scenarios including core-powered mass loss, photoevaporation, or that these planets are primordially rocky. Interpretations differ on the physical mechanism of atmospheric mass-loss, but the result is the same - primordially accreted atmospheres are removed in such a way that different planets are affected in different ways over different timescales, resulting in a valley which separates a population of stripped-core planets (super-Earths) from those which have retained their H/He envelopes (sub-Neptunes) (Owen & Wu, 2013; Ginzburg et al., 2018; Gupta & Schlichting, 2019). Whilst investigations have been launched into the existence of planets in the valley, it is still not immediately clear if there are any firm detections of planets in the gap. Larger population studies have indicated that at least a few planets fall within the radius valley, but conversely smaller and more precise planet samples have revealed a complete lack of planets across the gap (Ho & Eylen, 2023). It has been observed, however, that the valley varies with planetary orbital separation, system age, and mass and metallicity of the host stars, which is consistent with the predictions of atmospheric mass-loss models (David et al., 2021; Sandoval et al., 2021; Petigura et al., 2022).
The internal structure of sub-Neptunes is not just limited to that of a rocky core surrounded by a gaseous atmosphere, it has been theorised that these planets might hold significant fractions of ices or liquid water (Rogers & Seager, 2010; Dorn et al., 2017). Zeng et al. (2021) suggested that the radii of planets hotter than 900K and with masses below 20 M⊕ can be reproduced assuming ice-dominated compositions without significant gaseous envelopes. The entire mass-radius (M-R) distribution of planets around the radius valley could therefore be separated into three groups: purely rocky planets, gas-poor water worlds, and gas-rich water worlds, where the first two represent the two peaks in the radius distribution (Luque & Pallé, 2022). However, Rogers et al. (2023) argues that the existence of small planets with hydrogen atmospheres are consistent with the data, once thermal evolution and mass-loss are properly accounted for. This means that there is a strong degeneracy between water-world and silicate/iron-hydrogen models, and that the characterisation of larger sub-Neptunes that reside in this area of the M-R diagram can be used to determine planetary evolution and formation pathways (Kubyshkina & Fossati, 2022).
@@ 2.1.2 CHEOPS
The data were processed using the CHEOPS Data Reduction Pipeline (DRP v13; Hoyer et al. 2020) that conducts frame calibration, instrumental and environmental correction, and aperture photometry using pre-defined radii ( = 22.5″[RINF], 25.0″[DEFAULT], 30.0″[RSUP], and OPTIMAL radius) as well as a noise-optimised radius [ROPT]. The size of the OPTIMAL radius is determined on a per-visit basis to account for the differing brightnesses of different targets, and is determined by minimizing the signal-to-noise ratio (Hoyer et al., 2020). Both visits of HD 77946 have the same optimized aperture radius of 25 pixels. The DRP produced flux contamination (see Hoyer et al. 2020 and Wilson et al. 2022 for computation and usage) that was subtracted from the light curves. We retrieved the data and corresponding instrumental basis vectors, assessed the quality using the pycheops Python package (Maxted et al., 2022), and decorrelated with the parameters suggested by this package, which can be found in Table 1. Outliers were also trimmed from the light curves, with points that were 4 away from the median value removed. We concluded that using the OPTIMAL radius was the best option as it gave the highest signal-to-noise. Therefore, we used these detrended data, shown in Figure 2, for further analysis.
@@ 2.2.1 DRS
Following the confident detection of the transiting planet with the DRS RVs, the s-BART (Silva et al., 2022) template-matching pipeline was applied to the same data-set. This approach is built around the core assumption that the RV signal introduced by the presence of an orbiting planet should be achromatic, i.e. independent of the wavelength at which it is measured. Telluric features are accounted for through a binary mask that rejects any wavelength for which Earth’s transmittance drops under 1% of its continuum value, which is enlarged to consider the BERV motion. Furthermore, to ensure that each observation is informed by the same wavelength region, s-BART only uses the spectral orders that are common to all observations in the dataset. This model was implemented in a Bayesian framework, allowing for a straightforward and consistent characterization of the uncertainties associated with the RV extraction, and has been yielding good results for both ESPRESSO data (e.g. Faria et al. 2022) and HARPS-N (Wilson et al., 2024). This resulted in an average RV precision of 1.4 ms-1 and an RMS of 4.45 ms-1. The HARPS-N s-BART data are presented in Table 6.
@@ 3.1 Atmospheric parameters
We calculate the U, V, and W Galactic velocities using the method of Johnson & Soderblom (1987) and data from the most recent Gaia data release (for examples see Mortier et al. (2020); Wilson et al. (2022)), and compute the thin and thick disk, and halo membership probabilities. This was done using the framework outlined in Wilson et al. (2024), which combines new Gaia offset-corrected (Lindegren et al., 2021) data (Gaia Collaboration et al., 2023) with a known probability method (Reddy et al., 2006), then uses a novel Monte Carlo approach to compute a weighted average combination of probabilities using four sets of velocity standards (Bensby et al., 2003; Bensby et al., 2014; Reddy et al., 2006; Chen et al., 2021) corrected for the Local Standard of Rest (Koval et al., 2009). We find U kms-1, V kms-1, W kms-1, resulting in 98.5%, 1.5%, and 0% thin disk, thick disk, and halo probabilities, respectively, which is in line with the given metallicity for this type of star.
@@
| RA | 09:07:06.63 | Gaia DR3 |
| Declination | +46:40:20.73 | Gaia DR3 |
| [mas] | Gaia DR3 | |
@@
| Prot [d] | This work | |
| RV [m s-1] | This work | |
| [km s-1] | This work | |
| [km s-1] | This work | |
| [km s-1] | This work | |
@@ 3.2 Stellar mass, radius, and age
Stellar parameters of HD 77946 were determined using isochrones and stellar evolution tracks. For a detailed discussion on this see Mortier et al. (2020). In brief, spectroscopically determined values for the effective temperature and metallicity from ARES+MOOG and SPC were used as inputs to the isochrones python package (Morton, 2015), together with the Gaia Early Data Release 3 (Brown et al., 2021) parallax and the apparent magnitude in 8 bands, listed in Table 2. Two sets of stellar evolution models were used, the Dartmouth Stellar Evolution Database (Dotter et al., 2008) and the MESA Isochrones and Stellar Tracks (MIST - Dotter, 2016). For each stellar evolution model (Dartmouth/MIST) the analysis was run twice; once using each of the spectroscopic inputs (ARES+MOOG/SPC), totalling to 4 individual runs. Final values for the stellar mass, radius, and age were obtained by combining the posteriors from each run, and taking the median and the 16th and 84th percentile. We derived values M⊙, R⊙, and Gyr. Combining the derived mass and radius, we can compute a new value for the surface gravity, which is consistent with, but more precise than the spectroscopic surface gravity.
@@ 3.3 Stellar activity and rotation
Signals intrinsic to stars themselves still pose significant problems, as they can induce RV variations that can drown out or even mimic planetary signals (Rajpaul et al., 2021). Modelling and thus mitigating these signals is notoriously difficult. However, a technique that has proved successful for disentangling stellar activity signals from planetary signals is to use a Gaussian Process (GP) framework (Haywood et al., 2014; Rajpaul et al., 2015; Stock et al., 2020). GPs have a number of features that make them very well suited to the joint modelling of activity processes and dynamical (e.g. planetary) signals, giving probabilistic estimates of the model parameters, which leads to more precise and accurate mass estimates.
@@
where and are two times of observation, and the other terms are the variability period of the star (), the variability amplitude (), the periodic characteristic length (associated with the number of spots/spot regions on the surface of the star, ), and the non-periodic characteristic length (associated with the spot decay timescale, ). The GP was run with log-uniform priors on the amplitude 0 ln() 2.7 and active-region evolution timescale 24 40 d. These boundaries were selected based on the stellar type of HD 77946, an F dwarf, so we assume its active regions evolve on approximately the same timescale as the Sun. The SPC gives kms-1, and based on a stellar radius of 1.31 R⊙ this limits the stellar rotation period to a value of below 22 days. Based on this estimate and a few initial runs with uninformative priors, the GP priors on the stellar rotation period were set to be uniform with 12 19 d. The log harmonic complexity was controlled with a uniform prior -1 ln() 0. In short, we used a model with up to five unknown Keplerian signals to maximise exploration, and stellar activity decorrelation performed against the scalpels basis vectors, incorporated with a quasi-periodic GP regression.
@@ 4.1 Joint Transit & RV Modelling
| [M⊕] | Planet mass | |||
| [R⊕] | Planet radius | |||
@@
Inferred stellar and planetary parameters from the combined fit are shown in Table 3, alongside the priors implemented. The prior for is explained in Section 3.3, whilst the prior on stellar density comes from stellar mass and radius estimates in Section 3.2. The priors placed on the limb-darkening coefficients were calculated using interpolation of Table 25 from Claret (2017) for the TESS light curves and Table 8 from Claret (2021) for the CHEOPS light curves. The fit converges to a moderately eccentric planetary orbit, with a posterior on corresponding to a detection. Using the estimates for the stellar mass and radius obtained in Section 3.2 and the HARPS-N DRS RVs, we derive that HD 77946 b has a mass M⊕ and a radius R⊕.
We also performed the same combined GP, transit, and RV fit using the s-BART RVs, as mentioned in Section 2.2.2, with the resulting inferred stellar and planetary parameters shown in Table 3. In this case the fit converged to a slightly more eccentric planetary orbit than that of the fit with the DRS RVs, which led to a slightly higher posterior on corresponding to a detection. From this we also derived a slightly lower mass and radius for HD 77946 b of M⊕ and R⊕. These results still fall within of those derived from the HARPS-N DRS RVs, but this does suggest that the choice of data reduction technique can influence the resulting planetary parameters. We favour these results as we find that the fit with the s-Bart RVs gives a stronger detection of the planet’s mass and required less jitter in the fit, although there is tension in the derived eccentricity when using either RV dataset.
@@
We ensured that the nested sampling was converging to consistent results by running the same model multiple times, and across different machines. Given the aforementioned problems in constraining the stellar rotation period, we also ran models with different priors on , despite this the planetary parameter results were still consistent to 1 on all runs. A corner plot with all the planetary parameters can be found in Appendix A, see Figure 11. In addition to this the RV semi-amplitude (2.69 ± 0.57 ms-1) obtained from the scalpels analysis in Section 3.3 was combined with the stellar mass and orbital inclination (1.17 M⊙ and 87.10 degrees respectively) which lead to a mass determination of 8.83 ± 1.86 M⊕. This result is consistent with the results from the joint GP, photometric, and spectroscopic fit determined using PyORBIT. Fits were also performed with and without a GP, and comparing the Bayesian Information Criterion (BIC), the non-GP fit had a log-median BIC of -64134 whilst the fit with a GP had a log-median BIC of -56987. Trotta (2008) suggests under the assumption of a multivariate Gaussian posterior that a log-Bayes factor of should be considered "strong" evidence for a model. In this case the log-Bayes factor is 3573.5 (BIC), further justifying the use of a GP.
In order to further confirm this, a phase folded plot of the HARPS-N RVs colour coded to the time period they were taken was produced (see Appendix A Figure 10). From the grouping of RVs in this plot, we were able to conclude that the results were not being skewed by only sampling some parts of the orbital phase. An additional check was performed to ensure the validity of the eccentricity obtained, in which a 1-planet fit was attempted where the planets orbit was forced to be circular (with ). From this we calculate a slightly lower mass, M⊕, corresponding to a similar confidence detection of , with R⊕. Comparing the Bayesian Information Criterion (BIC), the forced circular orbit had a log-median BIC of -57009 whilst the eccentric orbit had a log-median BIC of -56987. Thus, whilst the confidence detections on are both around , the median BIC of the eccentric orbit is better. In addition to this, the log-Bayes factor is 11 (BIC), which supports choosing the more eccentric model. Orbits for single-planet systems are also not typically zero (Van Eylen et al., 2019) and there is no compelling argument to fix the eccentricity of this planet as circular, therefore we favour the model with an eccentric orbit.